### v1

# ars548_ros: An ARS 548 RDI radar driver for ROS2

Abstract
The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long range radar sensors with new RF antenna arrays, which offer digital beam forming. This radar measures independently the distance, speed and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation radarDesc . Unfortunately, there were not any drivers available for Linux systems to make the user able to analyze the data acquired from this sensor to the best of our knowledge. In this paper, we present a driver that is able to interpret the data from the ARS 548 RDI sensor and produce data in Robot Operation System version 2 (ROS2). Thus, this data can be stored, represented and analyzed by using the powerful tools offered by ROS2. Besides, our driver offers advanced object features provided by the sensor, such as relative estimated velocity and acceleration of each object, its orientation and angular velocity. We focus on the configuration of the sensor and the use of our driver and advanced filtering and representation tools, offering a video tutorial for these purposes. Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor for baseline purposes is available, so that the user can check the correctness of our driver.
keywords:
ROS2, ARS 548 RDI, Radar Sensor, FMPC radar
journal: SoftwareX
Code metadata
Current code version
1.0
Permanent link to code/repository used for this code version
https://github.com/robotics-upo/ars548_ros
Permanent link to Reproducible Capsule
Legal Code License
BSD 3-Clause License
Code versioning system used
Git
Software code languages, tools, and services used
C++, C.
Compilation requirements, operating environments & dependencies
Linux
If available Link to developer documentation/manual
https://github.com/robotics-upo/ars548_ros/blob/master/Readme.md
Support email for questions
Software metadata
Current software version
1.0
Permanent link to executables of this version
Permanent link to Reproducible Capsule
Legal Software License
List one of the approved licenses
Computing platforms/Operating Systems
Linux
Installation requirements
ROS2 Humble
Support email for questions
dalejo@us.es
1 Motivation and significance
Autonomous navigation in reduced visibility scenarios, i.e. scenarios with the presence of fog, rain, dust or smoke, is still a great challenge for robots, as the prevalent sensors used for obstacle detection in robotics, i.e. Light Detection and Ranging sensors (LiDAR), RGB and RGB-D cameras, can significantly reduce their precision or even become useless when operating under these conditions booker_jfr_2007 ; li_icra_2021 . Also, sensors onboard a smartphone can be used in robotic application or for indoor positioning thanks to applications such as GetSensorData GUTIERREZ2022101186 , but the performance of most onboard sensors would also drop in reduced visibility scenarios. Therefore, different sensors should be employed to allow autonomous robots to be deployed in reduced visibility scenarios.
One alternative can be the use of thermal infrared cameras for odometry and localization khattak2020keyframe ; saputra2020deeptio . In this paper, we focus on RAdio Detection And Ranging sensors (radar), as aerosols do not affect them because their wavelength is larger than most aerosol particles. Thus, radar sensors are able to penetrate rain, dust and smoke. However, using a larger wavelength also has drawbacks, as the measurements obtained by a radar sensor have worse distance and angular accuracy and resolution when compared to LiDAR measurements Mielle2019 .
Localization and navigation in reduced-visibility conditions is an active research field in autonomous car navigation. The vast majority of reduced-visibility datasets include RADAR information oxford_radar ; seeing_through_fog ; hong_radar_slam_2020 . A common approach is to perform data fusion of information perceived from LiDAR and RADAR sensors to ensure that the objects are detected in all circumstances. Unfortunately, the rotating radar sensors used in these datasets (such as the Navtech Radar CTS350-X) are too expensive and heavy to be included in a cost-effective robotic solution.
In the past years, advances in Frequency-Modulated Continuous Wave (FMCW) radar, which are based on the Doppler-effect to detect surrounding objects, have allowed the development of new radar sensors for robotic applications. Not only are these sensors able to detect the distance of nearby objects to them, but they can also estimate their relative velocity by taking advantage of Doppler effect. This velocity information can be used to improve the odometry estimation in harsh environments radar_odometry_quist_2016 , which is fundamental for the safe autonomous navigation of the robots. Still, their use for 3D navigation is challenging due to the low resolution and narrow vertical Field of View (FoV) of inexpensive models, such as the IWR6843 from Texas Instruments.
More recently, the ARS 548 RDI sensor from Continental (see Figure 1) has been developed, which use more advanced modulation techniques to better estimate the position and velocity of nearby objects. We believe that this sensor can be a good option for enabling a mobile robot or autonomous car to operate in reduced visibility scenarios, as it is a good compromise solution between inexpensive but feature lacking FMCW radar sensors and 2D rotating radars used for automotive purposes.
In this paper, we describe the architecture of our software solution, highlighting its interface to Robotic Operating System v2 ros2 (ROS2), which allows us to use the representation, logging and processing tools offered by its community. We illustrate the good performance of our driver by including data collected from real experiments with the data of a OS1-32 Lidar from Ouster as a baseline of the data. The driver is distributed in a free license in the hope that it will be helpful to the community.
Figure 1: Continental ARS 548 RDI high resolution RADAR sensor.
2 Software description
We developed our driver with two main objectives in mind. On the one hand, our driver generates standard ROS2 messages such as PointCloud2 that includes information of the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects (see Figure 2). By using standard messages, the user can take advantage of existing representation and processing tools available in ROS2. On the other hand, we are interested in providing the user with all the information generated by the radar without any processing. To this end, we have created the following custom ROS 2 messages: ObjectList, DetectionList and Setup (see Figure 3), which are a direct translation of the data structures send by the sensor as described in the ARS 548 reference manual ars548_doc .
Figure 2: Basic flow diagram of ARS 548 driver.
2.1 Software architecture
The main components of our driver are listed on Figure 3. There are two main ROS2 packages in it, with the following main functionalities:
  • 1.
    ARS548_driver. This folder contains the libraries, code and functions needed for the actual driver to work. In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS548_messages), which are provided to the user via ROS2 middleware. On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor.
  • 2.
    ARS548_messages. In this folder we define the custom messages that are sent by our driver. These messages follow the data structure of the messages generated by the radar, as defined in ars548_doc .
Figure 3: Driver Architecture Scheme.
All of the code of the driver is available at our repository on GitHub repo .
2.2 Software functionalities
This section highlights the main functionalities provided by ARS 548 driver.
2.2.1 Reception and translation of messages.
The proposed driver is able to establish a UDP connection with the sensor and translate the different information from the sensor to custom and standard ROS2 messages, making the necessary endian translations by means of templated C++ functions.
With this functionality we provide the user with all the information of ARS548 as emitted by the sensors with our custom messages. Additionally, we provide standard messages which are easier to visualize and process with standard ROS2 packages.
2.2.2 Visualization of the results.
By generating standard ROS2 messages such as PointCloud2 and PoseArray, the user can easily represent the objects in the standard ROS2 visualizer (RViz2 rviz ). To this end, we provide the user with custom visualization configurations that can be loaded in RViz for easily representing the measurements from the radar sensor (see Figure 3). You can see an example of representation of the radar measurements in Figure 4.
Figure 4: Example of object measurements captured by the ARS 548 RDI sensor, displayed in RViz.
2.2.3 Object filtering
Finally, our ARS 548 driver package provides the user with a customizable filter that is able to interpret our custom messages and emit filtered point clouds depending on the extended information generated by the sensor. As an example, see Section 3.2 in which we used a filter to only represent moving vehicles on the environment using the speed estimation of the sensor.
3 Illustrative examples
The driver presented in this paper has been experimentally tested in field experiments with the robotic ARCO platform arco (see Figure 5). In particular, we describe here two example scenarios. First, a traffic monitoring scenario, in which the velocity estimation offered by the radar sensor can be of great use to distinguish the road traffic from the background. Second, a typical robotic scenario, in which the robot should be capable of both generating a map of the environment and localizing itself on it, in a problem that is usually referred as Simultaneous Localization and Mapping (SLAM).
Figure 5: Our ARCO robotic platform equipped with both an ARS 548 RDI sensor and an OS1-32 LiDAR sensor. The necessary equipment for the experiments is pointed out on the Figure.
3.1 Experimental setup
Our ARCO robotic platform has been equipped with an ARS-548 sensor and an OS1-32 LiDAR sensor, which was used as a reference to confirm the validity of the measurements obtained by the radar sensor (see Figure 5). We used an MSI GF62 7RE Laptop with 32GB RAM and Intel Core i7 processor connected to the ARS 548 RDI sensor and the OS1-32 LiDAR sensor via Ethernet, with independent connections. For each experiment, we generated one or more ROSBag files with all the information from both sensors. In particular, with regards to ARS 548 RDI, we saved both our custom topics and the standard ones.
3.2 Experiment 1. Traffic monitoring
In this experiment, our ARCO robotic platform remained static in the proximity of a road. The objective here is to use the ARS 548 RDI sensor to monitor the road traffic nearby, taking advantage of its medium-long range (up to 300m) and its capability of measuring the speed of the detected objects. We use the object filtering feature of our driver obtaining an object point cloud that contains vehicles moving faster than 10 m/s, which are assumed as road traffic (see Figure 6).
Figure 6: Snapshot of Experiment 1. (left) Radar detections on blue, objects faster than 10 km/h on big white dots and LiDAR detections on multicolor. (right) Geolocalized radar measures represented over a map from OpensStreetMap. Red lines represent cumulative radar detections of two different vehicles over the map of the environment.
3.3 Experiment 2. SLAM
Work in progress.
4 Impact
This driver is published with a free software license in the hope of minimizing development efforts of all users of the ARS 548 RDI sensor that would like to implement ROS2-based applications with it. By doing so, the users can easily setup the driver and acquire its data with ROS2 interfaces and then focus on processing the data as needed for their own experiments, significantly speeding-up their research and development process.
Our main goal is to use this device for inspection robotics application carried out in reduced visibility environments, mainly due to weather phenomena such as rain, fog and snow or emergency situations involving smoke. These applications may include offshore oil&gas extraction plants inspection, sewer inspection and the inspection of transmission and electric power lines rs16020355 . By using this sensor, we can ensure that the inspection operations can be safely performed in all circumstances.
Another application field is autonomous driving. In fact, this is the main are of application of ARS 548 RDI devices. The idea here is to incorporate this sensor in an autonomous vehicle to provide it with detailed information of surrounding objects including the type of object or vehicle, and its relative velocity and acceleration. With this information, the user can design collision detection and resolution algorithms that guarantee the collision-free operation of the vehicle in spite of harsh weather or poor illumination circumstances.
5 Conclusions
In this project we describe a ROS2 driver to acquire and process all the data sent by the ARS 548 RDI radar, show some experiments monitoring the traffic to test its processing speed and precision and give examples where the use of a radar is necessary.
Acknowledgements
Authors would like to thank Fernando Amodeo for his invaluable help configuring the VLAN connection to the ARS 548 RDI sensor.

### v2

@@
-ars548_ros: An ARS 548 RDI radar driver for ROS2
+ars548_ros: An ARS 548 RDI radar driver for ROS
 Abstract
 The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long range radar sensors with new RF antenna arrays, which offer digital beam forming.
-This radar measures independently the distance, speed and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation radarDesc .
-Unfortunately, there were not any drivers available for Linux systems to make the user able to analyze the data acquired from this sensor to the best of our knowledge.
-In this paper, we present a driver that is able to interpret the data from the ARS 548 RDI sensor and produce data in Robot Operation System version 2 (ROS2).
-Thus, this data can be stored, represented and analyzed by using the powerful tools offered by ROS2.
+This radar measures independently the distance, speed and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation [1].
+Unfortunately, to the best of our knowledge, there are no open source drivers available for Linux systems to enable users to analyze the data acquired by the sensor.
+In this paper, we present a driver that can interpret the data from the ARS 548 RDI sensor and make it available over the Robot Operating System versions 1 and 2 (ROS and ROS2).
+Thus, these data can be stored, represented, and analyzed using the powerful tools offered by ROS.
 Besides, our driver offers advanced object features provided by the sensor, such as relative estimated velocity and acceleration of each object, its orientation and angular velocity.
-We focus on the configuration of the sensor and the use of our driver and advanced filtering and representation tools, offering a video tutorial for these purposes.
-Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor for baseline purposes is available, so that the user can check the correctness of our driver.
-keywords:
-ROS2, ARS 548 RDI, Radar Sensor, FMPC radar
-journal: SoftwareX
+We focus on the configuration of the sensor and the use of our driver including its filtering and representation tools.
+Besides, we offer a video tutorial to help in its configuration process.
+Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor, to have baseline measurements, is available, so that the user can check the correctness of our driver [2].
+keywords: ROS ; Linux driver ; Radar Sensor
 Code metadata
-Current code version
-1.0
-Permanent link to code/repository used for this code version
-https://github.com/robotics-upo/ars548_ros
+Current code version 1.0.0
+Permanent link to code/repository used for this code version https://github.com/robotics-upo/ars548_ros
 Permanent link to Reproducible Capsule
-Legal Code License
-BSD 3-Clause License
-Code versioning system used
-Git
-Software code languages, tools, and services used
-C++, C.
-Compilation requirements, operating environments & dependencies
-Linux
-If available Link to developer documentation/manual
-https://github.com/robotics-upo/ars548_ros/blob/master/Readme.md
-Support email for questions
+Legal Code License BSD 3-Clause License
+Code versioning system used Git
+Software code languages, tools, and services used C++, C.
+Compilation requirements, operating environments & dependencies Linux
+If available Link to developer documentation/manual https://github.com/robotics-upo/ars548_ros/blob/master/Readme.md
+Support email for questions dalejo@us.es
 Software metadata
-Current software version
-1.0
+Current software version 1.0.0
 Permanent link to executables of this version
 Permanent link to Reproducible Capsule
-Legal Software License
-List one of the approved licenses
-Computing platforms/Operating Systems
-Linux
-Installation requirements
-ROS2 Humble
-Support email for questions
-dalejo@us.es
+Legal Software License List one of the approved licenses
+Computing platforms/Operating Systems Linux
+Installation requirements ROS2 Humble or ROS Noetic
+Support email for questions dalejo@us.es
 1 Motivation and significance
@@
 scenarios with the presence of fog, rain, dust or smoke, is still a great challenge for robots, as the prevalent sensors used for obstacle detection in robotics, i.e.
-Light Detection and Ranging sensors (LiDAR), RGB and RGB-D cameras, can significantly reduce their precision or even become useless when operating under these conditions booker_jfr_2007 ; li_icra_2021 .
-Also, sensors onboard a smartphone can be used in robotic application or for indoor positioning thanks to applications such as GetSensorData GUTIERREZ2022101186 , but the performance of most onboard sensors would also drop in reduced visibility scenarios.
+Light Detection and Ranging sensors (LiDAR), RGB and RGB-D cameras, can significantly reduce their precision or become useless when operating under these conditions [3, 4].
+Also, sensors onboard a smartphone can be used in robotic applications thanks to applications such as GetSensorData [5], but the performance of most onboard sensors would also drop in reduced visibility scenarios.
 Therefore, different sensors should be employed to allow autonomous robots to be deployed in reduced visibility scenarios.
-One alternative can be the use of thermal infrared cameras for odometry and localization khattak2020keyframe ; saputra2020deeptio .
+One alternative can be the use of thermal infrared cameras for odometry and localization [6, 7].
 In this paper, we focus on RAdio Detection And Ranging sensors (radar), as aerosols do not affect them because their wavelength is larger than most aerosol particles.
 Thus, radar sensors are able to penetrate rain, dust and smoke.
-However, using a larger wavelength also has drawbacks, as the measurements obtained by a radar sensor have worse distance and angular accuracy and resolution when compared to LiDAR measurements Mielle2019 .
+However, using a larger wavelength also has drawbacks, as the measurements obtained by a radar sensor have worse distance and angular accuracy and resolution when compared to LiDAR measurements [8].
 Localization and navigation in reduced-visibility conditions is an active research field in autonomous car navigation.
-The vast majority of reduced-visibility datasets include RADAR information oxford_radar ; seeing_through_fog ; hong_radar_slam_2020 .
-A common approach is to perform data fusion of information perceived from LiDAR and RADAR sensors to ensure that the objects are detected in all circumstances.
+In fact, the vast majority of reduced-visibility datasets include radar information [9, 10, 11].
+A common approach is to perform data fusion [12] of information perceived from LiDAR and radar sensors to ensure that the objects are detected in all circumstances.
 Unfortunately, the rotating radar sensors used in these datasets (such as the Navtech Radar CTS350-X) are too expensive and heavy to be included in a cost-effective robotic solution.
@@
 Not only are these sensors able to detect the distance of nearby objects to them, but they can also estimate their relative velocity by taking advantage of Doppler effect.
-This velocity information can be used to improve the odometry estimation in harsh environments radar_odometry_quist_2016 , which is fundamental for the safe autonomous navigation of the robots.
+This velocity information can be used to improve the odometry estimation in harsh environments [13], which is fundamental for the safe autonomous navigation of robots.
 Still, their use for 3D navigation is challenging due to the low resolution and narrow vertical Field of View (FoV) of inexpensive models, such as the IWR6843 from Texas Instruments.
-More recently, the ARS 548 RDI sensor from Continental (see Figure 1) has been developed, which use more advanced modulation techniques to better estimate the position and velocity of nearby objects.
+More recently, the ARS 548 RDI sensor from Continental (see Figure 1) has been developed.
+It uses more advanced modulation techniques to better estimate the position and velocity of nearby objects.
 We believe that this sensor can be a good option for enabling a mobile robot or autonomous car to operate in reduced visibility scenarios, as it is a good compromise solution between inexpensive but feature lacking FMCW radar sensors and 2D rotating radars used for automotive purposes.
-In this paper, we describe the architecture of our software solution, highlighting its interface to Robotic Operating System v2 ros2 (ROS2), which allows us to use the representation, logging and processing tools offered by its community.
-We illustrate the good performance of our driver by including data collected from real experiments with the data of a OS1-32 Lidar from Ouster as a baseline of the data.
-The driver is distributed in a free license in the hope that it will be helpful to the community.
+Unfortunately, we did not find any open source driver that would allow us to acquire and process the data acquired from this sensor.
 Figure 1: Continental ARS 548 RDI high resolution radar sensor.
+In this paper, we describe the architecture of the ars548_ros package.
+We highlight its interface to both versions of the Robotic Operating System [14] (ROS), which are becoming the standard for robotic development at least for academic purposes, even though ROS2 is more industrial focused than its predecessor, ROS.
+To develop ROS and ROS2 interfaces is very convenient as it allows us to use its representation, logging and processing tools.
+Hereinafter, we will use ROS when we refer to both versions of ROS.
+We illustrate the good performance of our driver by including data collected from real experiments with measurements of a OS1-32 LiDAR sensor from Ouster as a baseline.
+Finally, the driver is distributed in the BSD 3-Clause License in the hope that it will be helpful to the community.
 2 Software description
 We developed our driver with two main objectives in mind.
-On the one hand, our driver generates standard ROS2 messages such as PointCloud2 that includes information of the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects (see Figure 2).
-By using standard messages, the user can take advantage of existing representation and processing tools available in ROS2.
+On the one hand, our driver generates standard ROS messages such as PointCloud2 that include information of the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects.
+By using standard messages, the user can take advantage of existing representation and processing tools available in ROS.
 On the other hand, we are interested in providing the user with all the information generated by the radar without any processing.
-To this end, we have created the following custom ROS 2 messages: ObjectList, DetectionList and Setup (see Figure 3), which are a direct translation of the data structures send by the sensor as described in the ARS 548 reference manual ars548_doc .
-Figure 2: Basic flow diagram of ARS 548 driver.
+To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures send by the sensor as described in the ARS 548 reference manual [15].
+The messages have been included in Appendix A for the sake of completeness.
+Finally, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3).
+Figure 2: a) Basic flow diagram of ARS 548 driver. b) Example application of the filter: detection of objects moving faster than a threshold velocity.
 2.1 Software architecture
-The main components of our driver are listed on Figure 3. There are two main ROS2 packages in it, with the following main functionalities:
-  • 1.
+Our solution consists of two ROS packages, with the following main functionalities:
   • ARS548_driver.
 This folder contains the libraries, code and functions needed for the actual driver to work.
-In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS548_messages), which are provided to the user via ROS2 middleware.
+In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS548_messages), which are provided to the user via ROS middleware.
 On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor.
-  • 2.
-    ARS548_messages.
-In this folder we define the custom messages that are sent by our driver.
-These messages follow the data structure of the messages generated by the radar, as defined in ars548_doc .
-Figure 3: Driver Architecture Scheme.
-All of the code of the driver is available at our repository on GitHub repo .
+  • ARS548_messages. In this folder we define the custom messages that are sent by our driver. These messages follow the data structure of the messages generated by the radar, as defined in [15].
+The full driver source code is publicly available in our GitHub repository [16]. Also, please refer to Appendix B for the additional contents associated with the paper.
 2.2 Software functionalities
@@
 2.2.1 Reception and translation of messages.
-The proposed driver is able to establish a UDP connection with the sensor and translate the different information from the sensor to custom and standard ROS2 messages, making the necessary endian translations by means of templated C++ functions.
-With this functionality we provide the user with all the information of ARS548 as emitted by the sensors with our custom messages.
-Additionally, we provide standard messages which are easier to visualize and process with standard ROS2 packages.
+The proposed driver is able to establish a UDP connection with the sensor and translate the different information from the sensor to custom and standard ROS messages, making the necessary endian translations by means of templatized C++ functions.
+With this functionality we provide the user with all the information from ARS548 as emitted by the sensors with our custom messages.
+Additionally, we provide standard messages which are easier to visualize and process with standard ROS packages.
 2.2.2 Visualization of the results.
-By generating standard ROS2 messages such as PointCloud2 and PoseArray, the user can easily represent the objects in the standard ROS2 visualizer (RViz2 rviz ).
-To this end, we provide the user with custom visualization configurations that can be loaded in RViz for easily representing the measurements from the radar sensor (see Figure 3).
+By generating the PointCloud2 and PoseArray standard ROS messages, the user can easily represent the objects in the standard ROS visualizer (RViz [17]).
+To this end, we provide the user with custom visualization configurations that can be loaded in RViz for easily representing the measurements from the radar sensor (see Figure 2a).
 You can see an example of representation of the radar measurements in Figure 3.
-Figure 4: Example of object measurements captured by the ARS 548 RDI sensor, displayed in RViz.
+Figure 3: Radar detection measurements captured by the ARS 548 RDI sensor in blue and OS1 liDAR measurements in red. Please note the alignment between LiDAR and radar measurements.
 2.2.3 Object filtering
-Finally, our ARS 548 driver package provides the user with a customizable filter that is able to interpret our custom messages and emit filtered point clouds depending on the extended information generated by the sensor.
-As an example, see Section 3.2 in which we used a filter to only represent moving vehicles on the environment using the speed estimation of the sensor.
+The ars548_ros package provides the user with a filter able to interpret our custom messages and to provide filtered point clouds depending on the extended information generated by the sensor.
+The available information is defined in the Object message described in Appendix A.
+The filter can easily be customized via inheritance just by overriding a method; please refer to our repository for more details.
+As an example, see Section 3.2 in which we used a filter to represent only moving vehicles in the environment using the speed estimation from the sensor.
 3 Illustrative examples
-The driver presented in this paper has been experimentally tested in field experiments with the robotic ARCO platform arco (see Figure 5).
+The driver presented in this paper has been experimentally tested in field experiments with the robotic ARCO platform [18] (see Figure 4).
 In particular, we describe here two example scenarios.
@@
 3.1 Experimental setup
-Our ARCO robotic platform has been equipped with an ARS-548 sensor and an OS1-32 LiDAR sensor, which was used as a reference to confirm the validity of the measurements obtained by the radar sensor (see Figure 5).
-We used an MSI GF62 7RE Laptop with 32GB RAM and Intel Core i7 processor connected to the ARS 548 RDI sensor and the OS1-32 LiDAR sensor via Ethernet, with independent connections.
+Our ARCO robotic platform has been equipped with an ARS 548 RDI sensor and an OS1-32 LiDAR sensor, which was used as a reference to confirm the validity of the measurements obtained by the radar sensor (see Figure 4).
+We used an MSI GF62 7RE Laptop with 32GB RAM and Intel Core i7 processor connected to both sensors via Ethernet, with independent connections.
 For each experiment, we generated one or more ROSBag files with all the information from both sensors.
 In particular, with regards to ARS 548 RDI, we saved both our custom topics and the standard ones.
+The dataset is available at [2].
 3.2 Experiment 1. Traffic monitoring
@@
 The objective here is to use the ARS 548 RDI sensor to monitor the road traffic nearby, taking advantage of its medium-long range (up to 300m) and its capability of measuring the speed of the detected objects.
-We use the object filtering feature of our driver obtaining an object point cloud that contains vehicles moving faster than 10 m/s, which are assumed as road traffic (see Figure 6).
+We use the object filtering feature of our driver (see Figure 2b) obtaining an object point cloud that contains vehicles moving faster than 10 km/h, which are assumed as road traffic (see Figure 5).
 Figure 5: Snapshot of Experiment 1.
-(left) Radar detections on blue, objects faster than 10 km/h on big white dots and LiDAR detections on multicolor.
-(right) Geolocalized radar measures represented over a map from OpensStreetMap.
+(left) Radar detections on blue, Radar objects on red, Radar object moving faster than 10 km/h on a green sphere and LiDAR measurements on white.
+(right) Geolocalized radar measurements represented over a map from OpenStreetMap [19].
 Red lines represent cumulative radar detections of two different vehicles over the map of the environment.
 3.3 Experiment 2. SLAM
-Work in progress.
+In the second experiment, our ARCO robotic platform navigates within an open environment while estimating its trajectory through the information received by the radar.
+Figure 3 illustrates the alignment between the measurements from our driver and the LiDAR, verifying its proper functioning.
+To validate the results of this experiment, we have employed the odometry generated by the LeGO-LOAM package [20] as a baseline, which relies on the LiDAR sensor for data acquisition (see Figure 6).
+The data acquired from both the radar and LiDAR sensors will be used for future odometry and SLAM techniques, ensuring accurate mapping and localization, especially in hazardous scenarios where radar is employed.
+Figure 6: The baseline trajectory generated by the LeGO-LOAM package is represented in large green dots over the generated map projection in 2D.
+The LiDAR measurements are represented in red dots, with small green dots representing features detected by Lego-LOAM.
 4 Impact
-This driver is published with a free software license in the hope of minimizing development efforts of all users of the ARS 548 RDI sensor that would like to implement ROS2-based applications with it.
-By doing so, the users can easily setup the driver and acquire its data with ROS2 interfaces and then focus on processing the data as needed for their own experiments, significantly speeding-up their research and development process.
+This open source driver has been published under the BSD 3-Clause License in the hope of minimizing the development efforts of all users of the ARS 548 RDI sensor who would like to implement ROS-based applications with it.
+By doing so, the users can easily setup the driver and acquire its data with ROS interfaces and then focus on processing the data as needed for their own experiments, significantly speeding up their research and development process.
 Our main goal is to use this device for inspection robotics application carried out in reduced visibility environments, mainly due to weather phenomena such as rain, fog and snow or emergency situations involving smoke.
-These applications may include offshore oil&gas extraction plants inspection, sewer inspection and the inspection of transmission and electric power lines rs16020355 .
-By using this sensor, we can ensure that the inspection operations can be safely performed in all circumstances.
+These applications may include the inspection of different facilities such as offshore oil&gas extraction plants, sewers and transmission and electric power lines [21].
+By using this sensor, we can ensure that inspection operations can be performed safely under all circumstances.
 Another application field is autonomous driving.
-In fact, this is the main are of application of ARS 548 RDI devices.
-The idea here is to incorporate this sensor in an autonomous vehicle to provide it with detailed information of surrounding objects including the type of object or vehicle, and its relative velocity and acceleration.
+In fact, this is the main area of application for the ARS 548 RDI devices.
+The idea here is to incorporate this sensor into an autonomous vehicle to provide it with detailed information of surrounding objects including the type of object or vehicle and its relative velocity and acceleration.
 With this information, the user can design collision detection and resolution algorithms that guarantee the collision-free operation of the vehicle in spite of harsh weather or poor illumination circumstances.
 5 Conclusions
-In this project we describe a ROS2 driver to acquire and process all the data sent by the ARS 548 RDI radar, show some experiments monitoring the traffic to test its processing speed and precision and give examples where the use of a radar is necessary.
+This paper described a ROS driver to configure and interface with the ARS 548 RDI radar sensor.
+The driver works on both ROS and ROS2 versions, maximizing its impact on the community.
+The driver has been tested in two different field experiments, showcasing the advanced information from the sensor and the filtering capabilities of our driver.
+Future work will focus on developing new radar odometry approaches, and LiDAR-radar data fusion mechanisms.
+Declaration of competing interest
+The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Data availability
+All the data used in this paper is openly available at [2].
 Acknowledgements
-Authors would like to thank Fernando Amodeo for his invaluable help configuring the VLAN connection to the ARS 548 RDI sensor.
+Authors would like to thank Fernando Amodeo for his invaluable help.
+This work has been supported by the grants INSERTION PID2021-127648OB-C31 and RATEC PDC2022-133643-C21.
+Both grants have been funded by the Spanish Ministry of Science, Innovation and Universities with grant number 10.13039/501100011033 and “European Union NextGenerationEU/PRTR”.
+Appendix A . Custom ARS 548 messages
+Below you can find the definition of the custom ROS messages in ars548_messages.
+Detection.msg
+1float32 f_azimuthangle #Unaligned Detection Azimuth Angle
+2float32 f_azimuthanglestd #Azimuth Angle Std
+3uint8 u_invalidflags #Detection Invalid Flags
+4float32 f_elevationangle #Unaligned Detection Elevation Angle
+5float32 f_elevationanglestd #Elevation Angle Std
+6float32 f_range #Detection Radial Distance
+7float32 f_rangestd #Radial Distance Std
+8float32 f_rangerate #Detection Radial Velocity
+9float32 f_rangeratestd #Radial Velocity Std
+10int8 s_rcs #Detection RCS
+11uint16 u_measurementid #Detection ID
+12uint8 u_positivepredictivevalue #Existence Probability
+13uint8 u_classification #Detection Classification
+14uint8 u_multitargetprobabilitym #Multi-Target Probability
+15uint16 u_objectid #Associated Object ID
+16uint8 u_ambiguityflag #Probability for resolved velocity ambiguity
+17uint16 u_sortindex #tbd
+DetectionList.msg
+1std_msgs/Header header
+2uint64 crc #Checksum (E2E Profile 7) (Reserved)
+3uint32 length #Len (E2E Profile 7) (Reserved)
+4uint32 sqc #SQC (E2E Profile 7) (Reserved)
+5uint32 dataid #Data ID (E2E Profile 7) (Reserved)
+6uint32 timestamp_nanoseconds #Timestamp Nanoseconds
+7uint32 timestamp_seconds #Timestamp Seconds
+8uint8 timestamp_syncstatus #Timestamp Sync Status
+9uint32 eventdataqualifier #Event Data Qualifier (unused)
+10uint8 extendedqualifier #Extended Qualifier (unused)
+11uint16 origin_invalidflags #Sensor Position Invalid flags (unused)
+12float32 origin_xpos #Sensor X Position with reference to rear axle
+13float32 origin_xstd #Sensor X Position STD (unused)
+14float32 origin_ypos #Sensor Y Position with reference to rear axle
+15float32 origin_ystd #Sensor Y Position STD (unused)
+16float32 origin_zpos #Sensor Z Position with reference to rear axle
+17float32 origin_zstd #Sensor Z Position STD (unused)
+18float32 origin_roll #Sensor Roll Angle (unused)
+19float32 origin_rollstd #Sensor Roll Angle STD (unused)
+20float32 origin_pitch #Sensor Pitch Angle with alignment correction
+21float32 origin_pitchstd #Sensor Pitch Angle STD
+22float32 origin_yaw #Sensor Yaw Angle with alignment correction
+23float32 origin_yawstd #Sensor Yaw Angle STD
+24uint8 list_invalidflags #Invalid flags (unused)
+25Detection[800] list_detections #Detection Array
+26float32 list_radveldomain_min #Ambiguity free Doppler velocity range Min
+27float32 list_radveldomain_max #Ambiguity free Doppler velocity range Max
+28uint32 list_numofdetections #Number of Detections
+29float32 aln_azimuthcorrection #Azimuth Alignment Correction
+30float32 aln_elevationcorrection #Elevation Alignment Correction
+31uint8 aln_status #Status of alignment
+Object.msg
+1uint16 u_statussensor #tbd
+2uint32 u_id #Unique ID of object
+3uint16 u_age #Age of object
+4uint8 u_statusmeasurement #Object Status
+5uint8 u_statusmovement #Object Movement Status
+6uint16 u_position_invalidflags #tbd
+7uint8 u_position_reference #Reference point position
+8float32 u_position_x #X Position of reference point
+9float32 u_position_x_std #X Position Std
+10float32 u_position_y #Y Position of reference point
+11float32 u_position_y_std #Y Position Std
+12float32 u_position_z #Z Position of reference point
+13float32 u_position_z_std #Z Position Std
+14float32 u_position_covariancexy #Covariance X Y
+15float32 u_position_orientation #Object Orientation
+16float32 u_position_orientation_std #Orientation Std
+17uint8 u_existence_invalidflags #unused
+18float32 u_existence_probability #Probability of Existence
+19float32 u_existence_ppv #unused
+20uint8 u_classification_car #Car Classification Probability
+21uint8 u_classification_truck #Truck Classification Probability
+22uint8 u_classification_motorcycle #Motorcycle Classification Probability
+23uint8 u_classification_bicycle #Bicycle Classification Probability
+24uint8 u_classification_pedestrian #Pedestrian Classification Probability
+25uint8 u_classification_animal #Animal Classification Probability
+26uint8 u_classification_hazard #Hazard Classification Probability
+27uint8 u_classification_unknown #Unknown Classification Probability
+28uint8 u_classification_overdrivable #unused
+29uint8 u_classification_underdrivable #unused
+30uint8 u_dynamics_absvel_invalidflags #unused
+31float32 f_dynamics_absvel_x #X Absolute Velocity
+32float32 f_dynamics_absvel_x_std #X Absolute Velocity Std
+33float32 f_dynamics_absvel_y #Y Absolute Velocity
+34float32 f_dynamics_absvel_y_std #Y Absolute Velocity Std
+35float32 f_dynamics_absvel_covariancexy #Covariance Absolute Velocity X Y
+36uint8 u_dynamics_relvel_invalidflags #unused
+37float32 f_dynamics_relvel_x #X Relative Velocity
+38float32 f_dynamics_relvel_x_std #X Relative Velocity Std
+39float32 f_dynamics_relvel_y #Y Relative Velocity
+40float32 f_dynamics_relvel_y_std #Y Relative Velocity Std
+41float32 f_dynamics_relvel_covariancexy #Covariance Relative Velocity X Y
+42uint8 u_dynamics_absaccel_invalidflags #unused
+43float32 f_dynamics_absaccel_x #X Absolute Acceleration
+44float32 f_dynamics_absaccel_x_std #X Absolute Acceleration Std
+45float32 f_dynamics_absaccel_y #Y Absolute Acceleration
+46float32 f_dynamics_absaccel_y_std #Y Absolute Acceleration Std
+47float32 f_dynamics_absaccel_covariancexy #Covariance Absolute Acceleration X Y
+48uint8 u_dynamics_relaccel_invalidflags #unused
+49float32 f_dynamics_relaccel_x #X Relative Acceleration
+50float32 f_dynamics_relaccel_x_std #X Relative Acceleration Std
+51float32 f_dynamics_relaccel_y #Y Relative Acceleration
+52float32 f_dynamics_relaccel_y_std #Y Relative Acceleration Std
+53float32 f_dynamics_relaccel_covariancexy #Covariance Relative Acceleration X Y
+54uint8 u_dynamics_orientation_invalidflags #unused
+55float32 u_dynamics_orientation_rate_mean #Object Orientation Rate
+56float32 u_dynamics_orientation_rate_std #Orientation Rate Std
+57uint32 u_shape_length_status #(unused) Shape Length Status
+58uint8 u_shape_length_edge_invalidflags #(unused) Invalid Flags Shape Length
+59float32 u_shape_length_edge_mean #Mean Shape Length
+60float32 u_shape_length_edge_std #(unused) Shape Length Std
+61uint32 u_shape_width_status #(unused) Shape Width Status
+62uint8 u_shape_width_edge_invalidflags #(unused) Invalid Flags Shape Width
+63float32 u_shape_width_edge_mean #Mean Shape Width
+64float32 u_shape_width_edge_std #(unused) Shape Width Std
+ObjectList.msg
+1std_msgs/Header header
+2uint64 crc #Checksum (E2E Profile 7) (Reserved)
+3uint32 length #Len (E2E Profile 7) (Reserved)
+4uint32 sqc #SQC (E2E Profile 7) (Reserved)
+5uint32 dataid #Data ID (E2E Profile 7) (Reserved)
+6uint32 timestamp_nanoseconds #Timestamp Nanoseconds
+7uint32 timestamp_seconds #Timestamp Seconds
+8uint8 timestamp_syncstatus #Timestamp Sync Status
+9uint32 eventdataqualifier #(unused) Event Data Qualifier
+10uint8 extendedqualifier #(unused) Extended Qualifier
+11uint8 objectlist_numofobjects #Number of Objects
+12Object[50] objectlist_objects #Object Array
+Status.msg
+1uint32 timestamp_nanoseconds #Timestamp Nanoseconds
+2uint32 timestamp_seconds #Timestamp Seconds
+3uint8 timestamp_syncstatus #Timestamp Sync Status
+4uint8 swversion_major #Software version (major)
+5uint8 swversion_minor #Software version (minor)
+6uint8 swversion_patch #Software version (patch)
+7float32 longitudinal #Longitudinal sensor position (AUTOSAR)
+8float32 lateral #Lateral sensor position (AUTOSAR)
+9float32 vertical #Vertical sensor position (AUTOSAR)
+10float32 yaw #Sensor yaw angle (AUTOSAR)
+11float32 pitch #Sensor pitch angle (AUTOSAR)
+12uint8 plugorientation #Orientation of plug
+13float32 length #Vehicle length
+14float32 width #Vehicle width
+15float32 height #Vehicle height
+16float32 wheelbase #Vehicle wheelbase
+17uint16 maximundistance #Maximum detection distance
+18uint8 frequencyslot #Center frequency
+19uint8 cycletime #Cycle time
+20uint8 timeslot #Cycle offset
+21uint8 hcc #Country code
+22uint8 powersave_standstill #Power saving in standstill
+23uint32 sensoripaddress_0 #Sensor IP address
+24uint32 sensoripaddress_1 #Reserved
+25uint8 configurationcounter #Counter that counts up if new configuration has been received and accepted
+26uint8 status_longitudinalvelocity #Signals if current VDY is OK or timed out
+27uint8 status_longitudinalacceleration #Signals if current VDY is OK or timed out
+28uint8 status_lateralacceleration #Signals if current VDY is OK or timed out
+29uint8 status_yawrate #Signals if current VDY is OK or timed out
+30uint8 status_steeringangle #Signals if current VDY is OK or timed out
+31uint8 status_drivingdirection #Signals if current VDY is OK or timed out
+32uint8 status_characteristicspeed #(Unused) Signals if current VDY is OK or timed out.
+33uint8 status_radarstatus #Signals if Radar Status is OK
+34uint8 status_voltagestatus # Bitfield to report under- and overvoltage errors
+35uint8 status_temperaturestatus #Bitfield to report under- and overtemperature errors
+36uint8 status_blockagestatus #Current blockage state and blockage self test state.
+Appendix B . Supplementary data
+The video tutorial for configuring the sensor and the driver can be found at [22]. The repository of the driver can be found at [16] and the acquired dataset can be found at [2].

### v3

@@
 Abstract
-The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long range radar sensors with new RF antenna arrays, which offer digital beam forming.
-This radar measures independently the distance, speed and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation [1].
-Unfortunately, to the best of our knowledge, there are no open source drivers available for Linux systems to enable users to analyze the data acquired by the sensor.
+The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long-range radar sensors with new RF antenna arrays, which offer digital beam forming.
+This radar measures independently the distance, speed, and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation [1].
+Unfortunately, to the best of our knowledge, there are no open-source drivers available for Linux systems to enable users to analyze the data acquired by the sensor.
 In this paper, we present a driver that can interpret the data from the ARS 548 RDI sensor and make it available over the Robot Operating System versions 1 and 2 (ROS and ROS2).
@@
 Code metadata
-Current code version 1.0.0
+Current code version 1.0.1
 Permanent link to code/repository used for this code version https://github.com/robotics-upo/ARS548_ros
@@
 Code versioning system used Git
-Software code languages, tools, and services used C++, C.
+Software code languages, tools, and services used C++
 Compilation requirements, operating environments & dependencies Linux
@@
 Software metadata
-Current software version 1.0.0
+Current software version 1.0.1
 Permanent link to executables of this version
@@
 Therefore, different sensors should be employed to allow autonomous robots to be deployed in reduced visibility scenarios.
-One alternative can be the use of thermal infrared cameras for odometry and localization [6, 7].
+One alternative can be the use of thermal infrared cameras for odometry and localization [6, 7], but the problem with cameras is the unknown scale ratio.
 In this paper, we focus on RAdio Detection And Ranging sensors (radar), as aerosols do not affect them because their wavelength is larger than most aerosol particles.
@@
 Unfortunately, the rotating radar sensors used in these datasets (such as the Navtech Radar CTS350-X) are too expensive and heavy to be included in a cost-effective robotic solution.
-In the past years, advances in Frequency-Modulated Continuous Wave (FMCW) radar, which are based on the Doppler-effect to detect surrounding objects, have allowed the development of new radar sensors for robotic applications.
-Not only are these sensors able to detect the distance of nearby objects to them, but they can also estimate their relative velocity by taking advantage of Doppler effect.
+In the past years, advances in Frequency-Modulated Continuous Wave (FMCW) radar, which are based on the Doppler effect to detect surrounding objects, have allowed the development of new radar sensors for robotic applications.
+Not only are these sensors able to detect the distance of nearby objects to them, but they can also estimate their relative velocity by taking advantage of the Doppler effect.
 This velocity information can be used to improve the odometry estimation in harsh environments [13], which is fundamental for the safe autonomous navigation of robots.
@@
 It uses more advanced modulation techniques to better estimate the position and velocity of nearby objects.
-We believe that this sensor can be a good option for enabling a mobile robot or autonomous car to operate in reduced visibility scenarios, as it is a good compromise solution between inexpensive but feature lacking FMCW radar sensors and 2D rotating radars used for automotive purposes.
-Unfortunately, we did not find any open source driver that would allow us to acquire and process the data acquired from this sensor.
+We believe that this sensor can be a good option for enabling a mobile robot or autonomous car to operate in reduced visibility scenarios [23, 24].
+This radar is a good compromise solution between inexpensive but feature-lacking FMCW radar sensors and 2D rotating radars used for automotive purposes.
+Unfortunately, we did not find any open-source driver that would allow us to acquire and process the data acquired from it.
 Figure 1: Continental ARS 548 RDI high resolution radar sensor.
 In this paper, we describe the architecture of the ARS 548_ros package.
-We highlight its interface to both versions of the Robotic Operating System [14] (ROS), which are becoming the standard for robotic development at least for academic purposes, even though ROS2 is more industrial focused than its predecessor, ROS.
-To develop ROS and ROS2 interfaces is very convenient as it allows us to use its representation, logging and processing tools.
+We highlight its interface to both versions of the Robotic Operating System [14] (ROS), which are becoming the standard for robotic development at least for academic purposes, even though ROS2 is more industrially focused than its predecessor, ROS.
+To develop ROS and ROS2 interfaces is very convenient as it allows us to use their representation, logging, and processing tools.
 Hereinafter, we will use ROS when we refer to both versions of ROS.
-We illustrate the good performance of our driver by including data collected from real experiments with measurements of a OS1-32 LiDAR sensor from Ouster as a baseline.
-Finally, the driver is distributed in the BSD 3-Clause License in the hope that it will be helpful to the community.
+We illustrate the good performance of our driver by including data collected from real experiments with measurements of an OS1-32 LiDAR sensor from Ouster as a baseline.
+Finally, the driver is distributed under the BSD 3-Clause License in the hope that it will be helpful to the community.
 2 Software description
 We developed our driver with two main objectives in mind.
-On the one hand, our driver generates standard ROS messages such as PointCloud2 that include information of the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects.
+On the one hand, our driver generates standard ROS messages such as PointCloud2 that include information on the position and relative velocity, if available, of the detected points and objects; and PoseArray that indicates the direction of the movement of the detected objects.
 By using standard messages, the user can take advantage of existing representation and processing tools available in ROS.
 On the other hand, we are interested in providing the user with all the information generated by the radar without any processing.
-To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures send by the sensor as described in the ARS 548 reference manual [15].
-The messages have been included in Appendix A for the sake of completeness.
+To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures sent by the sensor as described in the ARS 548 reference manual [15].
+The messages have been included in A for the sake of completeness.
 Finally, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3).
@@
 On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor.
+Finally, the radar_setup utility allows the user to change some parameters of the radar, such as its maximum detection range, its IP address, and many more.
   • ARS 548_messages. In this folder we define the custom messages that are sent by our driver. These messages follow the data structure of the messages generated by the radar, as defined in [15].
-The full driver source code is publicly available in our GitHub repository [16]. Also, please refer to Appendix B for the additional contents associated with the paper.
+The full driver source code is publicly available in our GitHub repository [16]. Please refer to B for the additional contents associated with the paper.
 2.2 Software functionalities
-This section highlights the main functionalities provided by ARS 548 driver.
+This section highlights the main functionalities provided by the ARS 548 driver.
 2.2.1 Reception and translation of messages.
 The proposed driver is able to establish a UDP connection with the sensor and translate the different information from the sensor to custom and standard ROS messages, making the necessary endian translations by means of templatized C++ functions.
-With this functionality we provide the user with all the information from ARS548 as emitted by the sensors with our custom messages.
-Additionally, we provide standard messages which are easier to visualize and process with standard ROS packages.
+With this functionality, we provide the user with all the information from ARS 548 as emitted by the sensors with our custom messages.
+Additionally, we provide standard messages, which are easier to visualize and process with standard ROS packages.
 2.2.2 Visualization of the results.
@@
 To this end, we provide the user with custom visualization configurations that can be loaded in RViz for easily representing the measurements from the radar sensor (see Figure 2a).
-You can see an example of representation of the radar measurements in Figure 3.
+You can see an example of the representation of the radar measurements in Figure 3.
 Figure 3: Radar detection measurements captured by the ARS 548 RDI sensor in blue and OS1 LiDAR measurements in red. Please note the alignment between LiDAR and radar measurements.
@@
 The ARS 548_ros package provides the user with a filter able to interpret our custom messages and to provide filtered point clouds depending on the extended information generated by the sensor.
-The available information is defined in the Object message described in Appendix A.
+The available information is defined in the Object message described in A.
 The filter can easily be customized via inheritance just by overriding a method; please refer to our repository for more details.
 As an example, see Section 3.3 in which we used a filter to represent only moving vehicles in the environment using the speed estimation from the sensor.
+2.2.4 Time Stamp Options
+We offer two different options to stamp the information received from the ARS 548:
+  • Keep the original stamp.
+Use this option if the ARS 548 device is synchronized with the local PC, which can be achieved by means of the gPTP protocol.
+To this end, the user PC has to be configured in master mode, so that the sensor can detect it [^1] Please refer to https://linuxptp.sourceforge.net/ for configuring it..
+  • Override stamp with local time.
+In case there is not synchronization between the RADAR sensor and the PC, for example if its Ethernet interface does not support it, we offer an option to override the received timestamp using the local reception time.
+However, there could be a loss of precision due to incorrect stamping that can be noticeable in high-speed navigation.
+2.2.5 Radar configuration
+We have included a basic utility that allows the user to change the configuration parameters of the ARS548.
+The configuration parameters of the device can be divided into the following categories.
+The reader is referred to the SensorConfiguration.msg file in A to see the detailed list of the configuration parameters:
+  1. 1.
+    Mounting position.
+This set of parameters allows the user to configure the mounting position of the device w.r.t.
+the front axle of the vehicle (lateral, longitudinal and vertical parameters).
+Besides, the orientation of the device (yaw and pitch angles) and the plug (left or right) can be specified.
+These parameters are used together with the vehicle parameters in the auto-alignment feature of the ARS 548.
+  2. 2.
+    Vehicle parameters. They allow the user to specify the main parameters of the car, which are width, length, height and wheel base.
+  3. 3.
+    Radar configuration.
+They change the mode of operation of the radar.
+You can configure its maximum detection distance (from 99 to 1500 meters), choose between three different frequency slots.
+They can also refer to internal configuration of the Radar, such as its IP address and power saving options, to name a few.
 3 Illustrative examples
@@
 First, a traffic monitoring scenario, in which the velocity estimation offered by the radar sensor can be of great use to distinguish the road traffic from the background.
-Second, a typical robotic scenario, in which the robot should be capable of both generating a map of the environment and localizing itself on it, in a problem that is usually referred as Simultaneous Localization and Mapping (SLAM).
+Second, a typical robotic scenario, in which the robot should be capable of both generating a map of the environment and localizing itself on it, in a problem that is usually referred to as Simultaneous Localization and Mapping (SLAM).
 Figure 4: Our ARCO robotic platform equipped with both an ARS 548 RDI sensor and an OS1-32 LiDAR sensor. The necessary equipment for the experiments is pointed out on the Figure.
@@
 Our ARCO robotic platform has been equipped with an ARS 548 RDI sensor and an OS1-32 LiDAR sensor, which was used as a reference to confirm the validity of the measurements obtained by the radar sensor (see Figure 4).
-We used an MSI GF62 7RE Laptop with 32GB RAM and Intel Core i7 processor connected to both sensors via Ethernet, with independent connections.
+We used an MSI GF62 7RE Laptop with 32GB RAM and an 11th generation Intel Core i7 processor.
+The laptop is connected to both sensors via Ethernet, with independent connections.
 For each experiment, we generated one or more ROSBag files with all the information from both sensors.
@@
 The dataset is available at [2].
+3.2 Driver Performance
+We have performed a thirty-minute long experiment, in which we used our driver visualized the data from the ARS 548 sensor in real time with RViz.
+Our goal is to check the usual load and memory usage using the Ubuntu built-in system monitor.
+During the experiment, the driver process used a mere 2% of the processing power of one core.
+Furthermore, its memory usage was stable at around 7 MB.
+Therefore, the driver has a very low impact on the system resources.
 3.3 Experiment 1. Traffic monitoring
@@
 The objective here is to use the ARS 548 RDI sensor to monitor the road traffic nearby, taking advantage of its medium-long range (up to 300 m) and its capability of measuring the speed of the detected objects.
-We use the object filtering feature of our driver (see Figure 2b) obtaining an object point cloud that contains vehicles moving faster than 10 km/h, which are assumed as road traffic (see Figure 5).
+We use the object filtering feature of our driver (see Figure 2b) obtaining an object point cloud that contains vehicles moving faster than 10 km/h, which are assumed to be road traffic (see Figure 5).
 Figure 5: Snapshot of Experiment 1.
-(left) Radar detections on blue, Radar objects on red, Radar object moving faster than 10 km/h on a green sphere and LiDAR measurements on white.
+(left) Radar detections in blue, Radar objects in red, Radar object moving faster than 10 km/h in a green sphere and LiDAR measurements in white.
 (right) Geolocalized radar measurements represented over a map from OpenStreetMap [19].
@@
 By doing so, the users can easily set up the driver and acquire its data with ROS interfaces and then focus on processing the data as needed for their own experiments, significantly speeding up their research and development process.
+The code is published in the GitHub platform [16].
 Our main goal is to use this device for inspection robotics application carried out in reduced visibility environments, mainly due to weather phenomena such as rain, fog and snow or emergency situations involving smoke.
@@
 The driver has been tested in two different field experiments, showcasing the advanced information from the sensor and the filtering capabilities of our driver.
-Future work will focus on developing new radar odometry approaches, and LiDAR-radar data fusion mechanisms.
+Future work will focus on implementing interfaces to allow the ROS modules from the user to provide the ARS 548 RADAR with information, such as estimated velocity and accelerations, to perform its provided auto-alignment function, which is adequate for automotive purposes.
+Finally, we are developing new radar odometry approaches, and LiDAR-radar data fusion mechanisms.
 Declaration of competing interest
@@
 Data availability
-All the data used in this paper is openly available at [2].
+All the experimental data presented in this paper are openly available at [2].
 Acknowledgements
@@
 Both grants have been funded by the Spanish Ministry of Science, Innovation and Universities with grant number 10.13039/501100011033 and “European Union NextGenerationEU/PRTR”.
-Appendix A . Custom ARS 548 messages
+Appendix A Custom ARS 548 messages
 Below you can find the definition of the custom ROS messages in ARS 548_messages.
@@
 36uint8 status_blockagestatus #Current blockage state and blockage self test state.
-Appendix B . Supplementary data
+SensorConfiguration.msg
+1float32 longitudinal #Longitudinal sensor position (AUTOSAR)
+2float32 lateral #Lateral sensor position (AUTOSAR)
+3float32 vertical #Vertical sensor position (AUTOSAR)
+4float32 yaw #Sensor yaw angle (AUTOSAR)
+5float32 pitch #Sensor pitch angle (AUTOSAR)
+6uint8 plugorientation #Orientation of plug
+7float32 length #Vehicle length
+8float32 width #Vehicle width
+9float32 height #Vehicle height
+10float32 wheelbase #Vehicle wheelbase
+11uint16 maximumdistance #Maximum detection distance
+12uint8 frequencslot #Center frequency (if MaximumDistance < 190 m only Mid can be selected)
+13uint8 cycletime #Cycle time
+14uint8 timeslot #Cycle offset
+15uint8 hcc #Country code
+16uint8 powersave_standstill #Power saving in standstill
+17uint32 sensoripaddress_0 #Sensor IP address
+18uint32 sensoripaddress_1 #Reserved (Configure as 2852025457 or 169.254.116.113)
+19uint8 newsensormounting #Flag if new sensor mounting position shall be configured
+20uint8 newvehicleparameters #Flag if new vehicle parameters position shall be configured
+21uint8 newradarparameters #Flag if new radar parameter shall be configured
+22uint8 newnetworkconfiguration # Flag if new IP address shall be configured
+Appendix B Supplementary data
 The video tutorial for configuring the sensor and the driver can be found at [22]. The repository of the driver can be found at [16] and the acquired dataset can be found at [2].

### v4

@@
 The ARS 548 RDI Radar is a premium model of the fifth generation of 77 GHz long-range radar sensors with new RF antenna arrays, which offer digital beamforming.
-This radar measures independently the distance, speed, and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation [1].
+This radar measures independently the distance, speed, and angle of objects without any reflectors in one measurement cycle based on Pulse Compression with New Frequency Modulation.
 Unfortunately, to the best of our knowledge, there are no open-source drivers available for Linux systems to enable users to analyze the data acquired by the sensor.
@@
 Besides, we offer a video tutorial to help in its configuration process.
-Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor, to have baseline measurements, is available, so that the user can check the correctness of our driver [2].
-keywords: ROS ; Linux driver ; Radar Sensor
+Finally, a dataset acquired with this sensor and an Ouster OS1-32 LiDAR sensor, to have baseline measurements, is available so that the user can check the correctness of our driver.
+keywords: ROS ; Linux driver ; Radar sensor ; ARS 548
 Code metadata
@@
 Permanent link to Reproducible Capsule
-Legal Software License List one of the approved licenses
+Legal Software License BSD 3-Clause License
 Computing platforms/Operating Systems Linux
@@
 To this end, we have created the following custom ROS messages: ObjectList, DetectionList and Status (see Figure 2a), which are a direct translation of the data structures sent by the sensor as described in the ARS 548 reference manual [15].
-The messages have been included in A for the sake of completeness.
-Finally, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3).
+The messages have been included in Appendix A for the sake of completeness.
+Moreover, we can use the advanced information available in the Objects structure to filter them, so that only the required objects are taken into consideration (see Figure 2b and Section 2.2.3).
+Finally, we include the definition of the SensorConfiguration message, which is used in the configuration tool that can change the parameters of the ARS 548 RDI radar (see Section 2.2.5).
 Figure 2: a) Basic flow diagram of ARS 548 driver. b) Example application of the filter: detection of objects moving faster than a threshold velocity.
@@
 This folder contains the libraries, code and functions needed for the actual driver to work.
-In particular, the ars_548 driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ARS 548_messages), which are provided to the user via ROS middleware.
-On the other hand, ars_548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor.
+In particular, the ars548_driver driver node receives the data from the radar via Ethernet link and transforms it into standard messages and custom messages (defined in ars548_messages), which are provided to the user via ROS middleware.
+On the other hand, ars548_filter enables the user to distinguish different types of objects thanks to the advanced functionalities of the sensor.
 Finally, the radar_setup utility allows the user to change some parameters of the radar, such as its maximum detection range, its IP address, and many more.
   • ars548_messages. In this folder we define the custom messages that are sent by our driver. These messages follow the data structure of the messages generated by the radar, as defined in [15].
-The full driver source code is publicly available in our GitHub repository [16]. Please refer to B for the additional contents associated with the paper.
+The full driver source code is publicly available in our GitHub repository [16]. Please refer to Appendix B for the additional contents associated with the paper.
 2.2 Software functionalities
@@
 The ars548_ros package provides the user with a filter able to interpret our custom messages and to provide filtered point clouds depending on the extended information generated by the sensor.
-The available information is defined in the Object message described in A.
+The available information is defined in the Object message (see Appendix A).
 The filter can easily be customized via inheritance just by overriding a method; please refer to our repository for more details.
@@
 The configuration parameters of the device can be divided into the following categories.
-The reader is referred to the SensorConfiguration.msg file in A to see the detailed list of the configuration parameters:
+The reader is referred to the SenorConfiguration.msg file in Appendix A to see the detailed list of the configuration parameters:
   1. 1.
@@
 They change the mode of operation of the radar.
-You can configure its maximum detection distance (from 99 to 1500 meters), choose between three different frequency slots.
+You can configure its maximum detection distance (from 99 to 1500 meters), and choose between three different frequency slots.
 They can also refer to internal configuration of the radar, such as its IP address and power saving options, to name a few.
@@
 3.2 Driver Performance
-We have performed a thirty-minute long experiment, in which we used our driver visualized the data from the ARS 548 sensor in real time with RViz.
+We have performed a thirty-minute long experiment, in which we used our driver and visualized the data from the ARS 548 sensor in real time with RViz.
 Our goal is to check the usual load and memory usage using the Ubuntu built-in system monitor.
@@
 Appendix A Custom ARS 548 messages
-Below you can find the definition of the custom ROS messages in ARS 548_messages.
+Below you can find the definition of the custom ROS messages defined in the ars548_messages ROS package.
 Detection.msg